OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? -----mph

OpenStudy (aum):

Let speed of car B be x Speed of car A is (x - 15) Distance traveled by car B in 1.5 h = x * 1.5 Distance traveled by car A in 2 h = (x - 15) * 2 These two distances are equal Therefore, 1.5x = 2(x - 15) 1.5x = 2x - 30 subtract 1,5x 0 = 0.5x - 30 add 30 30 = 0.5x divide by 0.5 30/0.5 = x x = 60

OpenStudy (anonymous):

thanks

OpenStudy (aum):

you are welcome.